OpenStudy (anonymous):

A point moves around the circle x^2 +y^2 = 9. when the point is at (square root of -3, square root of 6) its x coordinate is increasing at the rate of 20 units per second. How fast is the y coordinate changing at that instant?

OpenStudy (anonymous):

Find the rate of increase of the function first: \[x^2+y^2=9\\\frac{dx^2}{dt}+\frac{dy^2}{dt}=0\\2xx'+2yy'=0\\y'=\frac{x}{y}x' \]

OpenStudy (anonymous):

Then plug all knowns in ;)

OpenStudy (anonymous):

im still confused :/

OpenStudy (anonymous):

@sweett2639 Firstly do you know how to take derivative?

OpenStudy (baldymcgee6):

First label all our variables and rates.